# Provably Accurate Shapley Value Estimation via Leverage Score Sampling

### v1

@@ Abstract
Originally introduced in game theory, Shapley values have emerged as a central tool in explainable machine learning, where they are used to attribute model predictions to specific input features. However, computing Shapley values exactly is expensive: for a general model with $n$ features, $O(2^{n})$ model evaluations are necessary. To address this issue, approximation algorithms are widely used. One of the most popular is the Kernel SHAP algorithm, which is model agnostic and remarkably effective in practice. However, to the best of our knowledge, Kernel SHAP has no strong non-asymptotic complexity guarantees. We address this issue by introducing Leverage SHAP, a light-weight modification of Kernel SHAP that provides provably accurate Shapley value estimates with just $O(n\log n)$ model evaluations. Our approach takes advantage of a connection between Shapley value estimation and agnostic active learning by employing leverage score sampling, a powerful regression tool. Beyond theoretical guarantees, we show that Leverage SHAP consistently outperforms even the highly optimized implementation of Kernel SHAP available in the ubiquitous SHAP library [Lundberg & Lee, 2017].
@@ 1 Introduction
Shapley values also find other applications in machine learning, including in feature and data selection. For feature selection, the value function $v(S)$ is taken to be the model loss when when using the features in $S$ (Marcílio & Eler, 2020; Fryer et al., 2021). For data selection, $v(S)$ is taken to be the the loss when using the data observations in $S$ (Jia et al., 2019; Ghorbani & Zou, 2019).
@@ 1.2 Our Contributions
Despite its ubiquity in practice, to the best of our knowledge, no non-asymptotic theoretical accuracy guarantees are known for Kernel SHAP when implemented with $m<2^{n}-2$ row samples (which corresponds to $m$ evaluations of the value function, $v$). We address this issue by proposing Leverage SHAP, a light-weight modification of Kernel SHAP that 1) enjoys strong theoretical accuracy guarantees and 2) consistently outperforms Kernel SHAP in experiments.
Leverage SHAP begins with the observation that a nearly optimal solution to the regression problem in Equation 2 can be obtained by sampling just $\tilde{O}(n)$ rows with probability proportional to their statistical leverage scores, a natural measure for the “importance” or “uniqueness” of a matrix row (Sarlós, 2006; Rauhut & Ward, 2012; Hampton & Doostan, 2015; Cohen & Migliorati, 2017). [^2] Technically, this fact is known for unconstrained least square regression. Some additional work is needed to handle the linear constraint in Equation 2, but doing so is relatively straightforward. This fact immediately implies that, in principal, we should be able to provably approximate $\bm{\phi}$ with a nearly-linear number of value function evaluations (one for each sampled row).
@@
For any $\epsilon>0$ and constant $\delta>0$, the Leverage SHAP algorithm uses $m=O(n\log n+n/\epsilon)$ evaluations of $v$ in expectation [^3] Formally, if we include the dependence on $\delta$, we require $m=O(n\log(n/\delta)+n/(\epsilon\delta))$. and $O(mn^{2})$ additional runtime to return estimated Shapley values $\tilde{\bm{\phi}}$ satisfying $\langle\tilde{\bm{\phi}},\mathbf{1}\rangle=v([n])-v(\emptyset)$ and, with probability $1-\delta$,
@@
In words, Theorem 1.1 establishes that, with a near-linear number of function evaluations, we can compute approximate Shapley values whose objective value is close to the true Shapley values. We also require $O(mn^{2})$ additional runtime to solve the linearly constrained regression problem in Equation 3, although this cost can be reduced in practice using, e.g. iterative methods.
By leveraging the fact that $\mathbf{Z}$ is a well-conditioned matrix, the bound in Equation 5 also implies a bound on the average squared error, $\|\tilde{\bm{\phi}}-\bm{\phi}\|_{2}^{2}$, which is provided in Section 4.
Beyond our theoretical results, we also show that leverage score sampling can be naturally combined with paired sampling, without sacrificing theoretical guarantees. As for Kernel SHAP, doing so improves its performance in experiments. Moreover, we show that a natural “without replacement” version of leverage score sampling (which is even simpler than the method employed by Kernel SHAP) leads to additional accuracy improvements when $m$ is large. Overall, these simple optimizations lead to an algorithm that consistently outperforms the optimized version of Kernel SHAP available in the SHAP Library. We illustrate this point in Figure 1.
More extensive experiments are included in Section 5. We find that the improvement of Leverage SHAP over Kernel SHAP is especially substantial in settings where $n$ is large in comparison to $m$ and when we only have access to noisy estimates of the Shapley values. This is often the case in applications to explainable AI where, as discussion earlier, $v(S)$ is an expectation over fucntions involving random baselines, which is estimated via a finte sample.
@@ 1.3 Related Work
A natural alternative is to try selecting subsets in such a way that allows them to be reused across multiple players (Illés & Kerényi, 2019; Mitchell et al., 2022). However, since each term in the summation involves both $v(S)$ and $v(S\cup\{i\})$, it is difficult to achieve high levels of sample reuse when working with the summation definition. One option is to split the sum into two, and separately estimate $\sum_{S\subseteq[n]\setminus\{i\}}v(S\cup\{i\})/\binom{n-1}{|S|}$ and $\sum_{S\subseteq[n]\setminus\{i\}}v(S)/\binom{n-1}{|S|}$. This allows substantial subset reuse, but tends to perform poorly in practice due to high variance in the individual sums Wang & Jia (2023). By solving a global regression problem that determines the entire $\bm{\phi}$ vector, Kernel SHAP can be viewed as a more effective way of reusing subsets to obtain Shapley values for all players.
@@
For active learning of linear models with $n$ features under the $\ell_{2}$ norm, the $O(n\log n+n/\epsilon)$ sample complexity required by leverge score sampling is known to be near-optimal. More complicated sampling methods can achieve $O(n/\epsilon)$, which is optimal in the worst-case (Batson et al., 2012; Chen & Price, 2019). However, it is not clear how to apply such methods efficiently to an exponentially tall matrix, as we manage to do for leverage score sampling.
@@ 2 Background and Notation
As discussed, the Kernel SHAP method is based on an equaivalence between Shapley values and the solution of a contrained regression problem involving $\mathbf{Z}$ and $\mathbf{y}$. Formally, we have:
@@
For completeness, we provide a self-contained proof of Lemma 2.1 in Appendix F. The form in Equation 7 inspires the heuristic choice to sample sets with probabilities proportional to $w(\|\mathbf{z}\|_{1})$ in Kernel SHAP, as larger terms in the sum should intutively be sampled with higher probability to reduce variance of the estimate. However, as discussed, a more principaled way to approximately solve least squares regression problems via subsampling is to use probabilities proportional to the statistical leverage scores. Formally, these scores are defined as follows:
@@ 3 Leverage SHAP
Lemma 3.1 is proven in Appendix H. Based on the equivalent, we consider methods that construct a sampling matrix $\mathbf{S}$ and return:
@@ 3.1 Analytical Form of Leverage Scores
$=w(\|\mathbf{z}\|_{1})\|\mathbf{z}\|_{1}(n-\|\mathbf{z}\|_{1})=\binom{n}{\|\mathbf{z}\|_{1}}^{-1}.\qed$ (12)
@@ 3.2 Our Algorithm
Note that our procedure is different from the with-replacement Kernel SHAP procedure described in Section 1.1. Sampling without replacement requires more care to avoid iterating over the exponential number of pairs in $\mathcal{Z}$. Fortunately, we can exploit the fact that there are only $n-1$ different set sizes, and all sets of the same size have the same leverage score. This allows us to determine how many sets of a different size should be sampled (by drawing a binomial random variable), and we can then sample the required number of sets of each size uniformly at random. Ultimately, we can collect $m$ samples in $O(mn^{2})$ time. Details are deferred to Appendix G.
@@ 4 Theoretical Guarantees
We prove this guarantee in Appendix A. The proof modifies the standard analysis of leverage scores for active least squares regression, which requires two components: 1) a subspace embedding guarantee, proven with a matrix Chernoff bound applied to a sum of rank-$1$ random matrices, and 2) and an approximate matrix-multiplication guarantee, proven with a second moment analysis. We replace these steps with 1) a matrix Bernstein bound applied to a sum of rank-$2$ random matrices and 2) a block-approximate matrix multiplication guarantee.
@@
The proof of Corollary 4.1 appears in Appendix B. The statement of the corollary is in terms of a problem-specific parameter $\gamma$ which intuitively measures how well the optimal coefficients $\bm{\phi}$ can reach the target vector $\mathbf{b}$. In Appendix C, we explore how the performance varies with $\gamma$ in practice. We find (see e.g., Figures 5, 6, 7) that performance does erode as $\gamma$ increases for all regression-based algorithms, suggesting that $\gamma$ is not an artifact of our analysis.
@@ 5 Experiments
In the experiments, we evaluate Leverage SHAP and Kernel SHAP based on how closely they align with the ground truth Shapley values. We run our experiments on eight popular datasets from the SHAP library (Lundberg & Lee, 2017). We find that Leverage SHAP can even outperform the highly optimized Kernel SHAP in the SHAP library, especially for large $n$ or when access to the set function is noisy.
@@
Implementation Details. In order to compute the ground truth Shapley values for large values of $n$, we use a tree-based model for which we can compute the exact Shapley values efficiently using Tree SHAP (Lundberg & Lee, 2017). All of our code is written in Python and can be found in the supplementary material. (We will make a public repository after the review process.) We use the SHAP library for the Optimized Kernel SHAP implementation, Tree SHAP, and the datasets. We use XGBoost for training and evaluating trees, under the default parameters.
Additional experiments. So that we do not overcrowd the plots, we show performance on Kernel SHAP, the Optimized Kernel SHAP implementation, and Leverage SHAP in the main body. We also run all of the experiments with the ablated estimators in Appendix D. The ablation experiments (see e.g., Figure 9) suggest that Bernoulli sampling (sampling without replacement) improves Leverage SHAP for small $n$ and the paired sampling improves Leverage SHAP for all settings. Because it is more interpretable, we report the $\ell_{2}$-norm error metric to compare the estimated to the ground truth Shapley values. We also report the linear objective error for similar experiments in Appendix E.
Figure 3 plots the performance for Kernel SHAP, the Optimized Kernel SHAP, and Leverage SHAP as the number of samples varies (we set $m=5n,10n,20n,\ldots,160n$). For each estimator, the line is the median error and the shaded region encompasses the first and third quartile. [^4] We report the median and quartiles instead of the mean and standard deviation since the mean minus standard deviation could be negative for the small error values in our experiments. Kernel SHAP gives the highest error in all settings of $m$ and $n$, pointing to the importance of the paired sampling and sampling without replacement optimizations in both Optimized Kernel SHAP and Leverage SHAP. Because both methods sample without replacement, Optimized Kernel SHAP and Leverage SHAP achieve essentially machine precision as $m$ approaches $2^{n}$. For all other settings, Leverage SHAP reliably outperforms Optimized Kernel SHAP; the third quartile of Leverage SHAP is generally at the median of Optimized Kernel SHAP. The analogous experiment for the linear objective error depicted in Figure 11 indicates the same findings.
@@
Mean 0.00261 0.0208 15.4 0.000139 0.00298 0.00324 0.0358 130.0
1st Quartile 5.69e-07 0.0031 3.71 1.48e-05 0.00166 0.00163 0.0106 33.5
2nd Quartile 9.52e-06 0.0103 8.19 3.86e-05 0.00249 0.00254 0.0221 53.6
3rd Quartile 0.00181 0.029 20.1 0.000145 0.00354 0.00436 0.0418 132.0
@@
Mean 3.28e-14 0.00248 2.33 1.81e-05 0.000739 0.000649 0.00551 21.8
1st Quartile 2.12e-14 0.000279 0.549 2.16e-06 0.00027 0.000187 0.000707 5.85
2nd Quartile 3.55e-14 0.00138 1.26 5.43e-06 0.000546 0.000385 0.0024 13.0
3rd Quartile 4.22e-14 0.0036 3.03 1.63e-05 0.00101 0.000964 0.00665 25.1
@@
Mean 3.28e-14 0.000186 0.63 5.21e-06 0.000458 0.000359 0.00385 14.7
1st Quartile 2.12e-14 1.91e-05 0.0631 6.3e-07 0.000139 9.51e-05 0.000333 3.6
2nd Quartile 3.55e-14 8.31e-05 0.328 2.33e-06 0.000376 0.000235 0.00149 8.9
3rd Quartile 4.22e-14 0.000231 0.769 7.09e-06 0.000617 0.000556 0.00401 15.3
@@
Beyond sample size, we evaluate how the estimators perform with noisy access to the set functions. This setting is particularly relevant for Shapley values in explainable AI, since the set function may be an expectation that is approximated in practice. Figure 4 shows plots performance as noise is added to the set functions: Instead of observing $v(S)$, the estimators observe $v(S)+\zeta$ where $\zeta\sim\mathcal{N}(0,\sigma^{2})$ is normally distributed with standard deviation $\sigma$ (we set $\sigma=0,5\times 10^{-3},1\times 10^{-2},5\times 10^{-2},\ldots,1$). As in the sample size experiments, Leverage SHAP meets or exceeds the other estimators, suggesting the utility of the algorithm as a robust estimator in explainable AI. The analogous experiment for the linear objective error depicted in Figure 12 indicates the same findings.

### v2

@@ Abstract
Originally introduced in game theory, Shapley values have emerged as a central tool in explainable machine learning, where they are used to attribute model predictions to specific input features. However, computing Shapley values exactly is expensive: for a general model with $n$ features, $O(2^{n})$ model evaluations are necessary. To address this issue, approximation algorithms are widely used. One of the most popular is the Kernel SHAP algorithm, which is model agnostic and remarkably effective in practice. However, to the best of our knowledge, Kernel SHAP has no strong non-asymptotic complexity guarantees. We address this issue by introducing Leverage SHAP, a lightweight modification of Kernel SHAP that provides provably accurate Shapley value estimates with just $O(n\log n)$ model evaluations. Our approach takes advantage of a connection between Shapley value estimation and agnostic active learning by employing leverage score sampling, a powerful regression tool. Beyond theoretical guarantees, we show that Leverage SHAP consistently outperforms even the highly optimized implementation of Kernel SHAP available in the ubiquitous SHAP library [Lundberg & Lee, 2017].
@@ 1 Introduction
Shapley values also find other applications in machine learning, including in feature and data selection. For feature selection, the value function $v(S)$ is taken to be the model loss when when using the features in $S$ (Marcílio & Eler, 2020; Fryer et al., 2021). For data selection, $v(S)$ is taken to be the loss when using the data observations in $S$ (Jia et al., 2019; Ghorbani & Zou, 2019).
@@ 1.2 Our Contributions
Despite its ubiquity in practice, to the best of our knowledge, no non-asymptotic theoretical accuracy guarantees are known for Kernel SHAP when implemented with $m<2^{n}-2$ row samples (which corresponds to $m$ evaluations of the value function, $v$). We address this issue by proposing Leverage SHAP, a lightweight modification of Kernel SHAP that 1) enjoys strong theoretical accuracy guarantees and 2) consistently outperforms Kernel SHAP in experiments.
Leverage SHAP begins with the observation that a nearly optimal solution to the regression problem in Equation 2 can be obtained by sampling just $\tilde{O}(n)$ rows with probability proportional to their statistical leverage scores, a natural measure for the “importance” or “uniqueness” of a matrix row (Sarlós, 2006; Rauhut & Ward, 2012; Hampton & Doostan, 2015; Cohen & Migliorati, 2017). [^2] Technically, this fact is known for unconstrained least square regression. Some additional work is needed to handle the linear constraint in Equation 2, but doing so is relatively straightforward. This fact immediately implies that, in principle, we should be able to provably approximate $\bm{\phi}$ with a nearly-linear number of value function evaluations (one for each sampled row).
@@
For any $\epsilon>0$ and constant $\delta>0$, the Leverage SHAP algorithm uses $m~{}=~{}O(n\log(\frac{n}{\delta})+n\frac{1}{\epsilon\delta})$ evaluations of $v$ in expectation and $O(mn^{2})$ additional runtime to return estimated Shapley values $\tilde{\bm{\phi}}$ satisfying $\langle\tilde{\bm{\phi}},\mathbf{1}\rangle=v([n])-v(\emptyset)$ and, with probability $1-\delta$,
@@
In words, Theorem 1.1 establishes that, with a near-linear number of function evaluations, we can compute approximate Shapley values whose objective value is close to the true Shapley values. We also require $O(mn^{2})$ additional runtime to solve the linearly constrained regression problem in Equation 3, although this cost can be reduced in practice using, e.g., iterative methods.
By leveraging the fact that $\mathbf{Z}$ is a well-conditioned matrix, the bound in Equation 5 also implies a bound on the average squared error, $\|\tilde{\bm{\phi}}-\bm{\phi}\|_{2}^{2}$, which is provided in Section 4. (In Appendix C, we discuss why estimating Shapley values with multiplicative error is NP-hard.)
Beyond our theoretical results, we also show that leverage score sampling can be naturally combined with paired sampling, without sacrificing theoretical guarantees. Moreover, we show that a natural “without replacement” version of leverage score sampling leads to additional accuracy improvements when $m$ is large. Overall, these simple optimizations lead to an algorithm that consistently outperforms the optimized version of Kernel SHAP available in the SHAP Library (this ‘Optimized Kernel SHAP’ algorithm also uses paired sampling and sampling without replacement). We illustrate this point in Figure 1.
More extensive experiments are included in Section 5. We find that the improvement of Leverage SHAP over Kernel SHAP is especially substantial in settings where $n$ is large in comparison to $m$ and when we only have access to noisy estimates of the Shapley values. This is often the case in applications to explainable AI where, as discussion earlier, $v(S)$ is an expectation over functions involving random baselines and estimated via a finite sample.
@@ 1.3 Related Work
A natural alternative is to try selecting subsets in such a way that allows them to be reused across multiple players (Illés & Kerényi, 2019; Mitchell et al., 2022). However, since each term in the summation involves both $v(S)$ and $v(S\cup\{i\})$, it is difficult to achieve high levels of sample reuse when working with the summation definition. One option is to split the sum into two, and separately estimate $\sum_{S\subseteq[n]\setminus\{i\}}v(S\cup\{i\})/\binom{n-1}{|S|}$ and $\sum_{S\subseteq[n]\setminus\{i\}}v(S)/\binom{n-1}{|S|}$. This allows substantial subset reuse, but tends to perform poorly in practice due to high variance in the individual sums (Wang & Jia, 2023). By solving a global regression problem that determines the entire $\bm{\phi}$ vector, Kernel SHAP can be viewed as a more effective way of reusing subsets to obtain Shapley values for all players.
@@
For active learning of linear models with $n$ features under the $\ell_{2}$ norm, the $O(n\log n+n/\epsilon)$ sample complexity required by leverage score sampling is known to be near-optimal. More complicated sampling methods can achieve $O(n/\epsilon)$, which is optimal in the worst-case (Batson et al., 2012; Chen & Price, 2019). However, it is not clear how to apply such methods efficiently to an exponentially tall matrix, as we manage to do for leverage score sampling.
@@ 2 Background and Notation
As discussed, the Kernel SHAP method is based on an equivalence between Shapley values and the solution of a constrained regression problem involving $\mathbf{Z}$ and $\mathbf{y}$. Formally, we have:
@@
For completeness, we provide a self-contained proof of Lemma 2.1 in Appendix H. The form in Equation 7 inspires the heuristic choice to sample sets with probabilities proportional to $w(\|\mathbf{z}\|_{1})$ in Kernel SHAP, as larger terms in the sum should intuitively be sampled with higher probability to reduce variance of the estimate. However, as discussed, a more principled way to approximately solve least squares regression problems via subsampling is to use probabilities proportional to the statistical leverage scores. Formally, these scores are defined as follows:
@@ 3 Leverage SHAP
When there are multiple $\mathbf{x}$ that minimize the objective $\left\|\mathbf{Ax-b}\right\|_{2}^{2}$, we define $\operatorname*{arg\,min}_{\mathbf{x}}\left\|\mathbf{Ax-b}\right\|_{2}^{2}$ as the $\mathbf{x}$ that also minimizes $\|\mathbf{x}\|_{2}^{2}$.
Lemma 3.1 is proven in Appendix J. Using the equivalent unconstrained regressio n problem, we consider methods that construct a sampling matrix $\mathbf{S}$ and return:
@@ 3.1 Analytical Form of Leverage Scores
$=w(\|\mathbf{z}\|_{1})(n\|\mathbf{z}\|_{1}-\|\mathbf{z}\|_{1}^{2})=w(\|\mathbf{z}\|_{1})\|\mathbf{z}\|_{1}(n-\|\mathbf{z}\|_{1})=\binom{n}{\|\mathbf{z}\|_{1}}^{-1}.\qed$ (12)
@@ 3.2 Our Algorithm
Note that our procedure is different from the with-replacement Kernel SHAP procedure described in Section 1.1. Sampling without replacement requires more care to avoid iterating over the exponential number of pairs in $\mathcal{Z}$. Fortunately, we can exploit the fact that there are only $n-1$ different set sizes, and all sets of the same size have the same leverage score. This allows us to determine how many sets of a different size should be sampled (by drawing a binomial random variable), and we can then sample the required number of sets of each size uniformly at random. Ultimately, we can collect $m$ samples in $O(mn^{2})$ time. Details are deferred to Appendix I.
@@ 4 Theoretical Guarantees
We prove this guarantee in Appendix A. The proof modifies the standard analysis of leverage scores for active least squares regression, which requires two components: 1) a subspace embedding guarantee, proven with a matrix Chernoff bound applied to a sum of rank-$1$ random matrices, and 2) an approximate matrix-multiplication guarantee, proven with a second moment analysis. We replace these steps with 1) a matrix Bernstein bound applied to a sum of rank-$2$ random matrices and 2) a block-approximate matrix multiplication guarantee.
@@
The proof of Corollary 4.1 appears in Appendix B. The statement of the corollary is in terms of a problem-specific parameter $\gamma$ which intuitively measures how well the optimal coefficients $\bm{\phi}$ can reach the target vector $\mathbf{b}$. In Appendix E, we explore how the performance varies with $\gamma$ in practice. We find (see e.g., Figures 5, 6, 7) that performance does erode as $\gamma$ increases for all regression-based algorithms, suggesting that $\gamma$ is not an artifact of our analysis.
@@ 5 Experiments
In the experiments, we evaluate Leverage SHAP and Kernel SHAP based on how closely they align with the ground truth Shapley values. We primarily use the $\ell_{2}$-norm error i.e., $\|\bm{\phi}-\tilde{\bm{\phi}}\|_{2}^{2}/\|\bm{\phi}\|_{2}^{2}$ as the $\ell_{2}$ norm error. We run our experiments on eight popular datasets from the SHAP library (Lundberg & Lee, 2017). We find that Leverage SHAP can even outperform the highly optimized Kernel SHAP in the SHAP library, especially for large $n$ or when access to the set function is noisy.
@@
Implementation Details. In order to compute the ground truth Shapley values for large values of $n$, we use a tree-based model for which we can compute the exact Shapley values efficiently using Tree SHAP (Lundberg & Lee, 2017). All of our code is written in Python and can be found on Github333github.com/rtealwitter/leverageshap. We use the SHAP library for the Optimized Kernel SHAP implementation, Tree SHAP, and the datasets. We use XGBoost for training and evaluating trees, under the default parameters.
Additional experiments. So that we do not overcrowd the plots, we show performance on Kernel SHAP, the Optimized Kernel SHAP implementation, and Leverage SHAP in the main body. We also run all of the experiments with the ablated estimators in Appendix F. The ablation experiments (see e.g., Figure 9) suggest that Bernoulli sampling (sampling without replacement) improves Leverage SHAP for small $n$ and the paired sampling improves Leverage SHAP for all settings. Because it is more interpretable, we report the $\ell_{2}$-norm error metric to compare the estimated to the ground truth Shapley values. We also report the linear objective error for similar experiments in Appendix G.
Figure 3 plots the performance for Kernel SHAP, the Optimized Kernel SHAP, and Leverage SHAP as the number of samples varies (we set $m=5n,10n,20n,\ldots,160n$). For each estimator, the line is the median error and the shaded region encompasses the first and third quartile over 100 random runs. [^4] We report the median and quartiles instead of the mean and standard deviation since the mean minus standard deviation is negative for the small error values in our experiments. Kernel SHAP gives the highest error in all settings of $m$ and $n$, pointing to the importance of the paired sampling and sampling without replacement optimizations in both Optimized Kernel SHAP and Leverage SHAP. Because both methods sample without replacement, Optimized Kernel SHAP and Leverage SHAP achieve essentially machine precision as $m$ approaches $2^{n}$. For all other settings, Leverage SHAP reliably outperforms Optimized Kernel SHAP; the third quartile of Leverage SHAP is generally at the median of Optimized Kernel SHAP. The analogous experiment for the linear objective error depicted in Figure 11 indicates the same findings.
@@
Mean 0.026 0.0266 0.0553 0.0673 0.0465 0.0264 0.0604 0.12
1st Quartile 1.61e-05 0.00829 0.0116 0.0182 0.0244 0.0134 0.0202 0.0563
2nd Quartile 0.000898 0.0236 0.0229 0.0345 0.0404 0.0217 0.0388 0.089
3rd Quartile 0.0328 0.0424 0.0524 0.0936 0.056 0.0303 0.0843 0.149
@@
Mean 4.84e-09 0.00342 0.0093 0.00989 0.0117 0.00474 0.00758 0.0233
1st Quartile 1.66e-13 0.000802 0.00161 0.00187 0.00499 0.00194 0.00156 0.00962
2nd Quartile 2.17e-13 0.00238 0.00356 0.00489 0.00916 0.00391 0.00425 0.0173
3rd Quartile 2.69e-10 0.00489 0.00868 0.0122 0.015 0.00695 0.00871 0.0325
@@
Mean 4.84e-09 0.000311 0.0023 0.00477 0.00716 0.00288 0.00532 0.0156
1st Quartile 1.66e-13 4.47e-05 0.000215 0.000477 0.00289 0.000843 0.000995 0.0062
2nd Quartile 2.17e-13 0.000133 0.000969 0.00124 0.00528 0.00257 0.00288 0.0104
3rd Quartile 2.69e-10 0.000366 0.00241 0.00354 0.00891 0.00417 0.00554 0.0225
@@
Beyond sample size, we evaluate how the estimators perform with noisy access to the set functions in Appendix D. As in the sample size experiments, Leverage SHAP meets or exceeds the other estimators, suggesting the utility of the algorithm as a robust estimator in explainable AI.